# SoK: “Interoperability vs Security” Arguments: A Technical Framework

### v1

@@ Abstract
Concerns about big tech’s monopoly power have featured prominently in recent media and policy discourse, and regulators across the US, the EU, and beyond have ramped up efforts to promote healthier competition in the market. One of the favored approaches is to require certain kinds of interoperation between platforms, to mitigate the current concentration of power in the biggest companies. Unsurprisingly, interoperability initiatives have generally been met with vocal resistance by big tech companies. Perhaps more surprisingly, a significant part of that pushback has been in the name of security—that is, arguing against interoperation on the basis that it will undermine security.
We conduct a detailed examination of “security vs. interoperability” arguments in the context of recent antitrust proceedings in the US and the EU. First, we propose a taxonomy of such arguments. Second, we provide several detailed case studies, which illustrate our taxonomy’s utility in disentangling where security and interoperability are and are not in tension, where securing interoperable systems presents novel engineering challenges, and where “security arguments” against interoperability are really more about anti-competitive behavior than security. Third, we undertake a comparative analysis that highlights key considerations around the interplay of economic incentives, market power, and security across diverse contexts where security and interoperability may appear to be in tension. We believe systematically distinguishing cases and patterns within our taxonomy and analytical framework can be a valuable analytical tool for experts and non-experts alike in today’s fast-paced regulatory landscape.
@@ 1 Introduction
Billions of users of phones and other digital devices today entrust vast amounts of personal data to their devices and to services that are controlled, in general, by a small number of very large companies. These companies have had huge discretion over whom to share this data with, and how to secure these vast data flows—in messaging, digital payments, tap-and-go technology, and more. Recent concerns around concentration of power in big tech have led to calls and legal mandates for platforms to support increased interoperation. The idea is that this would open the market to more smaller players, promoting healthier competition, and break away from existing large-scale business models involving expansive control over software and hardware ecosystems by a single market actor. Large tech platforms have every incentive to resist moves to open up their systems. In this context, companies seeking to resist interoperability mandates are increasingly arguing that the security of their systems requires keeping them closed and limiting competition. They also portray their systems’ security as a key business proposition, and a key reason consumers are attracted to their business models.
@@
Such arguments are often framed as technical arguments about incurring unacceptable security risks, that courts and policymakers should be deferent to. These seemingly technical arguments may carry particular weight—for an audience of policymakers, regulators, and the general public—when invoked by companies that are already responsible for vast amounts of personal data that could be put at risk by an ill-advised regulation, especially those companies which have an established reputation for security and privacy.
Yet the “security vs. interoperability” argument is an inherently interdisciplinary one, about integrating technical security goals with other policy goals. For non-technical stakeholders, it is not easy to get a detailed sense of the technical considerations underlying the stated security concerns. For the security community, it can be challenging to get a detailed sense of the socio-technical, economic, and policy considerations essential to a holistic evaluation of the security considerations in the relevant regulatory context. Focusing on only technical aspects or only economic or regulatory aspects yields an inherently incomplete picture; policymaking based on this incomplete picture could lead to missteps in the important and fast-moving area of big tech and antitrust.
Thus, we believe it critical for the security community to engage with these questions not by opining on isolated technical questions—a natural approach to staying within one’s expertise—but by embracing the broader scope of the questions beyond security, especially when industry interests may not align with highlighting this broader scope.
This paper provides a framework for doing just that. Our framework captures the key economic considerations relevant to the regulatory context in which the “security vs. interoperability” discourse is rapidly unfolding. The framework is designed to be usable without detailed knowledge of antitrust law or the economics considerations in any given deployment context, and has sufficient generality to capture key aspects and patterns across all of the real-world case studies we present in this paper, which span diverse areas of security. Through analysis of these case studies, we illustrate our framework’s utility in disentangling where security and interoperability are and are not in tension, where securing interoperable systems presents novel engineering challenges, and where “security arguments” against interoperability may be a distraction from the real issue at hand: fair competition.
@@
The concentration of economic and political power in the hands of a few technology companies, alongside the novelty, complexity, and opacity of modern technology, has contributed to a climate in which companies strategically deploy (seemingly) technical arguments as devices to further their political and economic power and avoid regulation. For example, recent legal scholarship has highlighted the strategic use of privacy, content moderation, and security arguments to further companies’ economic and policy interests [45, 52, 34, 30, 50]. Notably, [45] discusses how privacy is used inconsistently as a pretext to further business aims, including not to allow competitors access to data, by companies that engage in expansive data selling and sharing of data when beneficial for business.
@@
The defining characteristic of such arguments is not that they are necessarily incorrect about security (or privacy), but that they appear to instrumentalize “security” into a label that is convenient for business ends. Although companies should, of course, be able to put forward the best case for their own interests in such regulatory contexts, this trend is concerning if and when seemingly technical arguments are treated with more deference than business or policy arguments, as more objective or factual, or as matters on which courts and regulators should defer to companies’ expertise.
This paper focuses in particular on the security-related arguments that companies like Apple raise in the context of recent antitrust investigations such as the Epic v. Apple case [^1] Epic Games, Inc. v. Apple Inc., Case 4:20-cv-05640-YGR, Findings of Fact and Conclusions of Law Proposed by Epic Games, Inc. 2021. in the US or proceedings to enforce the European Union’s new Digital Markets Act [^2] REGULATION (EU) 2022/1925 on contestable and fair markets in the digital sector (Digital Markets Act). (DMA) reform. In this context, tech companies raise a range of security-related justifications as a shield to avoid complying with competition and interoperability mandates that lawyers, courts and regulators are seeking to impose [39, 38, 59, 42, 22].
@@
When companies raise security arguments in this context, in general, the criticality and nuances of the security issue cannot be expected to be transparent to lawyers, courts and policymakers, or to other public stakeholders. Further confusing matters, some of the very same companies have a history of well-founded warnings against serious security risks in other policy contexts (such as backdooring encryption [26]). [^3] These latter security-critical arguments may well have served the companies’ business interests as well.
@@
In legal and regulatory proceedings, companies seeking to defend themselves from competition claims may strategically bring in security experts; then, those seeking to impose competition obligations will, in turn, bring their own security experts. In the Epic v. Apple case, companies’ experts have asserted that a company’s systems might suffer serious security risks or degradation if the company is forced to comply with interoperability and competition obligations, and the opposing experts have argued that this is not true [48].
Asked to adjudicate apparently conflicting technical evidence, courts, lawyers, and policymakers are often ill equipped to interrogate the nuances of the security arguments and trade-offs, to (efficiently) distinguish how critical a security argument is, and to determine which security arguments are more motivated by economic or policy ends than security goals. This can cause delays and raise costs whether in judicial or regulatory proceedings or legislative reform.
The risks to society and to the integrity of the legal and policy process can be significant. Big tech companies can hire very large teams of lawyers, lobbyists and security experts. Their strategic use of security arguments in service of their business goals can be misleading and difficult to sort through. It can significantly slow down judicial and regulatory processes, and lead to bad decisions or policies that undermine competition, openness and innovation, harming competitors, small businesses and consumers. Such decisions can, in turn, have long term effects on the structure of digital markets, on the diversity of players in them and on the quality of digital services available to consumers.
@@
We believe this paper’s focus on systematic evaluation of companies’ security arguments both in terms of technical security considerations and in the context of the broader economic and policy context of these proceedings can help the security and policy communities better analyze and conceptualize “security vs interoperability” tensions (and the lack thereof), leading to better policy decisions that will shape the future of technology markets and ecosystems.
@@
That security engineering requires consideration of contextual sociotechnical factors is hardly a new observation [2], and the security economics literature has long highlighted and studied the idea that “security failure is caused by bad incentives at least as often as by bad design” [3, 1]. Much of security economics considers how incentives influence security outcomes. Our inquiry takes a different angle, in a similar spirit: we ask how security arguments can be leveraged to serve other economic goals.
Digital competition policy and antitrust enforcement have been responding to digital markets trends long before the recent wave of interest in big tech and antitrust. Microsoft’s bundling of Internet Explorer with Windows was the subject of a landmark 1998 antitrust case. [^4] United States v. Microsoft Corporation, 253 F.3d 34 (D.C. Cir. 2001). Net neutrality regulations, which required internet service providers not to give preferred treatment to certain types of content or traffic for economic reasons, have been at the heart of political controversies since the 1990s [40, 60, 56, 61, 62, 57, 36]. The “security vs. interoperability” narrative, our focus in this paper, is a more recent development that very little prior work has considered. Since the DMA, a promising initial literature has begun to examine the complex technical considerations around securely interoperating end-to-end encrypted messaging [23, 43].
@@
This paper critically examines recent and ongoing “security vs. interoperability” arguments in the context of recent US and EU antitrust proceedings, including detailed technical analysis of the nature of claimed tensions between security and interoperability in diverse contexts. Our three key contributions are as follows.
@@
    We systematize “security vs. interoperability” arguments into a taxonomy consisting of two broad categories—engineering and vetting concerns—and a third hybrid category, and explain their relationship to the economic concepts of vertical and horizontal interoperability (§2).
@@
    We present seven case studies across diverse contexts in which “security vs. interoperability” arguments arise: messaging, app sideloading, in-app purchases, app stores, browser engines, NFC for host card emulation, and physical device interoperation (§3).
@@
    By analyzing these case studies through the lens of our framework, we illuminate key considerations around the interplay of economic incentives, market power, and security in “security vs. interoperability” arguments, and highlight the essential role of security trade-offs and existing policies in critically evaluating such arguments in their broader economic context (§§34).
2 Our Framework

### v2

# SoK: “ Security vs. Interoperability” Arguments: A Technical Framework

@@ Abstract
Concerns about big tech’s monopoly power have featured prominently in recent media and policy discourse, as regulators across the EU, the US, and beyond have ramped up efforts to promote healthier market competition. One favored approach is to require certain kinds of interoperation between platforms, to mitigate the current concentration of power in the biggest companies. Unsurprisingly, interoperability initiatives have generally been met with resistance by big tech companies. Perhaps more surprisingly, a significant part of that pushback has been in the name of security—that is, arguing against interoperation on the basis that it will undermine security.
We conduct a systematic examination of “security vs. interoperability” (SvI) discourse in the context of EU antitrust proceedings. Our resulting contributions are threefold. First, we propose a taxonomy of SvI concerns in three categories: engineering, vetting, and hybrid. Second, we present an analytical framework for assessing real-world SvI concerns, and illustrate its utility by analyzing several case studies spanning our three taxonomy categories. Third, we undertake a comparative analysis that highlights key considerations around the interplay of economic incentives, market power, and security across our diverse case study contexts, identifying common patterns in each taxonomy category. Our contributions provide valuable analytical tools for experts and non-experts alike to critically assess SvI discourse in today’s fast-paced regulatory landscape.
@@ 1 Introduction
Billions of users of phones and other digital devices today entrust vast amounts of personal data to their devices and to services that are controlled, in general, by a small number of very large companies. These companies have huge discretion over whom to share this data with, and how to secure these vast data flows—in messaging, digital payments, tap-and-go technology, and more.
Recent concerns around concentration of power in big tech have led to calls and legal mandates for platforms to support increased interoperation [45, 43, 72, 48, 23]. The idea is that this would open the market to more smaller players, promoting healthier competition, and break away from existing business models involving expansive control over software and hardware ecosystems by a single market actor. Large tech platforms have every incentive to resist moves to open up their systems. In this context, companies seeking to resist interoperability mandates are increasingly arguing that the security of their systems requires keeping them closed and limiting competition. They portray their systems’ security as a key business proposition, and a key reason consumers are attracted to their business models.
@@
Such arguments are often framed as technical arguments about incurring unacceptable security risks—an area outside the usual expertise of courts and policymakers, and squarely within that of the tech companies subject to regulation. [^1] Not only because of technological expertise in general, but also because of companies’ detailed knowledge of how their own systems work. These seemingly technical arguments may carry particular weight—for an audience of policymakers, regulators, and the general public—when invoked by companies that are already responsible for vast amounts of personal data that could be imperiled by ill-advised regulation, especially those which have established reputations for security and privacy.
Yet the “security vs. interoperability” narrative is an inherently interdisciplinary one, about acceptable tradeoffs in integrating technical security goals with other policy goals. Focusing on only technical aspects or only economic or regulatory aspects yields an inherently incomplete picture—policymaking based on which could lead to missteps in the important and fast-moving area of big tech and antitrust. Thus, we believe it critical for the security community to engage with these questions not by opining on isolated technical questions—a natural approach to staying within one’s expertise—but by embracing the broader scope of the questions beyond security, especially when industry interests may not align with highlighting this broader scope.
This paper provides a blueprint for doing just that. Our taxonomy divides “security vs. interoperability” tensions into three key categories: engineering, vetting, and hybrid. Our analytical framework, illustrated via a series of detailed case studies, captures the key economic considerations relevant to the regulatory context in which the “security vs. interoperability” discourse is rapidly unfolding. The framework is designed to be usable without knowledge of antitrust law or economics, and captures key aspects and patterns across all of the real-world cases we systematically examine. For each case study, we provide detailed analysis of the security, interoperability, and incentives considerations, which illustrate our approach’s utility in disentangling where security and interoperability are and are not in tension, where securing interoperable systems presents novel engineering challenges, and where “security arguments” against interoperability may be a distraction from the real issue: fair competition. Finally, our comparative analysis identifies common patterns in the interplay of incentives, market power, and security in each of our taxonomy categories.
@@
The concentration of economic and political power in the hands of a few technology companies, alongside the novelty, complexity, and opacity of modern technology, has contributed to a climate in which companies may strategically deploy (seemingly) technical arguments to further their political and economic power and avoid regulation. Recent legal scholarship has highlighted companies’ strategic use of privacy, content moderation, and security arguments to further economic and policy interests [53, 62, 37, 30, 60]. van Loo [53] discusses how companies use privacy inconsistently as a pretext to further business aims, including to limit competitors’ access to data, while engaging widely in data selling and sharing when beneficial for business.
@@
The defining characteristic of such arguments is not that they are necessarily incorrect about security (or privacy), but that they may instrumentalize “security” into a label that is convenient for business ends. Although companies should be able to put forward the best case for their own interests in such regulatory contexts, this trend is concerning if and when seemingly technical arguments are treated with more deference than business or policy arguments, as more objective or factual, or as matters on which courts and regulators should defer to companies’ expertise.
@@
This paper focuses on companies’ security-related discussion in the context of antitrust investigations in the European Union that involve mandated interoperability. When companies raise security arguments in this context, in general, the criticality and nuances of the security issue cannot be expected to be transparent to lawyers, courts and policymakers, or to other public stakeholders. Further confusing matters, some of the very same companies have a history of well-founded warnings against serious security risks in other policy contexts (such as backdooring encryption [33]). [^2] These latter security-critical arguments may well have served the companies’ business interests as well.
@@
Companies seeking to defend themselves from competition claims may strategically bring in security experts. Those seeking to impose competition obligations will, in turn, bring their own security experts. In the Epic v. Apple case, companies’ experts have asserted that a system might suffer serious security risks or degradation if the company is forced to comply with interoperability obligations, and the opposing experts have argued that this is not true [56].
@@
We believe this paper’s focus on systematic evaluation of companies’ security discussion can help the security and policy communities better analyze and conceptualize “security vs interoperability” tensions (and the lack thereof), leading to better policy decisions that will shape the future of technology markets and ecosystems.
@@
That security engineering requires consideration of contextual sociotechnical factors is hardly a new observation [2]. The security economics literature has long highlighted and studied the idea that “security failure is caused by bad incentives at least as often as by bad design” [3, 1]. Much of security economics considers how incentives influence security outcomes. Our inquiry takes a different angle, in a similar spirit: we ask how security arguments can be leveraged to serve other economic goals.
Digital competition policy and antitrust enforcement have been responding to digital markets trends long before the recent wave of interest in big tech and antitrust. Microsoft’s bundling of Internet Explorer with Windows was the subject of a landmark 1998 antitrust case. [^3] United States v. Microsoft Corporation, 253 F.3d 34 (D.C. Cir. 2001). Net neutrality regulations, which required internet service providers not to give preferred treatment to certain types of content or traffic for economic reasons, have been at the heart of political controversies since the 1990s [46, 73, 69, 74, 75, 70, 39]. The “security vs. interoperability” narrative, our focus in this paper, is a more recent development that very little prior work has considered. Since the European Union’s Digital Markets Act (DMA), a promising initial literature has begun to examine the complexities of securely interoperating end-to-end encrypted messaging [24, 49].
@@
We critically examine “security vs. interoperability” (SvI) discourse related to antitrust proceedings in the European Union (EU). Our four key contributions are as follows.
@@
    We conduct a systematic review of EU competition cases involving interoperability mandates for digital technology, which feature SvI considerations (§2).
@@
    We propose a three-part taxonomy of SvI concerns, and explain the taxonomy’s relationship to the economic concepts of vertical and horizontal interoperability (§3).
@@
    We introduce an analytical framework centered around four critical questions, and apply it to six case studies across diverse contexts featuring SvI concerns: messaging, app sideloading, app stores, browser engines, NFC, and physical device interoperation (§4).
  4. 4.
    We offer a comparative analysis of these case studies through the lens of our framework, illuminating key patterns in the interplay of economic incentives, market power, and security in each taxonomy category (§§45).
2 Our Methodology
We systematically reviewed cases where security comes up in the context of mandated interoperability in antitrust proceedings. We focus on EU cases, as there has been extensive and well documented surge of relevant antitrust enforcement activity in the EU in the past few years. Similar issues have arisen in US antitrust proceedings; occasionally we will note examples of prominent parallel US cases.
We first collected all cases in which there have been interoperability mandates related to digital technology, in the EU Commission’s competition case database [34] (henceforth, ‘EU Database’). We then searched for discussions about security in the context of our selected cases. The resulting collection of discourse—from legal and regulatory proceedings, press releases and press coverage, developer-facing documentation, and consumer-facing information—forms the basis for our systematization and analysis.
The rest of this section details: (1) how we selected our cases; (2) high-level results of our screening process; (3) how we tagged these cases with relevant technologies; and (4) how we identified relevant security issues.
What we mean by ‘interoperation’
We use interoperation to refer broadly to practices in which products or services produced by one company (e.g., smartphones) support or facilitate user engagement with products or services developed by a third-party (e.g., independent apps). For our purposes, interoperation comprises both technical interoperability (e.g., what are the requirements for an app to be technically capable of running on a given platform?) and interoperation policy (e.g., which technically interoperable apps does a platform permit as a matter of policy?).
2.1 Search Methodology
The EU Database features five categories of ‘policy areas,’ of which only three are within our scope: ‘Antitrust & Cartels,’ ‘Digital Markets Act,’ and ‘Merger.’ [^4] The other categories are ‘Foreign Subsidies’ and ‘State Aid.’
In our initial search, we searched all cases in either the ‘Antitrust & Cartels’ or the ‘Digital Markets Act’ policy area. In a follow-up search, we covered ‘Mergers’ with narrowed search criteria (including limited economic activities), a refinement we discuss below. In both the initial and the follow-up search, we used the search terms ‘interoperate interoperability interoperation,’ which the EU Database search engine interprets as a disjunction of the three terms.
Given the specific definition of interoperation we seek to capture, we conducted a preliminary search of DMA cases in the EU Database to check whether our search terms were used with the meaning we intended to capture. We found that the search terms were used in reference to both technical interoperability and interoperation policy, as defined above. We also tried other similar terms such as ‘compatible’ or ‘integrated’ and found these did not capture any cases we deemed relevant that were not in our original search.
From the results of our initial and follow-up searches, we manually screened for cases satisfying the following additional criteria: (1) the case concerned interoperation of digital technologies (not, e.g., trains), and (2) the case related to interoperation mandated by antitrust authorities. Section 2.2 offers more detail on the screening process.
Based on our initial search of Antitrust and DMA cases, we manually identified which classifications of economic activities [^5] These are classifications of industry made by the EU [36] included digital technologies relevant to our project scope. We used these classifications [^6] We included the following codes: NACE Rev 2 (2008): C.26.1, C.26.2, C.26.3, C.26.4, C.26.51, J.61.2, J.61.9, J.62, J.63, K.64.99; NACE Rev 2.1 (2025): C.26.1, C.26.2, C.26.3, C.26.4, C.26.51, K, L.64.9. to filter and improve efficiency in our follow-up search of Merger cases. [^7] We first eliminated codes that were facially unlikely to be about digital technology (e.g., C.10.52 - Manufacture of ice cream), and included any codes directly pertaining to digital technology. For the remaining codes (e.g., those pertaining to finance), we omitted codes associated only with one or more non-relevant cases in the initial search, and kept all others.
2.2 Screening for Relevant Cases
Our screening process consisted of two main phases: an initial screening based on a quick read, and an second phase where a more thorough reading was necessary.
Phase 1 (P1)
We filtered out any cases that were dismissed or that were just DMA ‘gatekeeper’ designations as these cannot give rise directly to interoperability mandates. We moreover filtered out cases that were quickly identifiable as out of scope, for one of the following reasons. First, many cases used the term interoperability only once, in a common definition of ‘Standards’—we discarded cases where this was the only mention of the term. Second, some cases mentioned interoperability a few times only as an existing product feature or part of the regulatory context; these cases did not involve mandates in their mentions of interoperability, so we discarded them. Third, sometimes multiple cases appear in the same document; we discarded cases in the same document as a mention of interoperability, that do not themselves mention the term. Phase 1 eliminated the majority of irrelevant cases (see Table I).
Phase 2 (P2)
All cases that remained required a more thorough reading to establish whether they were within our scoping criteria. We removed a number of cases that discussed interoperability as a product feature or important context in greater depth than was readily apparent, but did not mention interoperability mandates. We also removed one case that was not about digital technology.
Policy area Initial count Removed in P1 Removed in P2 Kept
Antitrust 28 22 4 2
DMA 17 13 0 4
Merger 66 45 13 8
TABLE I: Number of cases by policy area
The resulting 14 cases are listed in Table IV (Appendix).
Note
Among the Merger cases, many discussed potential interoperability issues but did not ultimately have an interoperability mandate. Generally this was because the Commission decided that the potential benefit to the merged entity of degrading interoperability was either not incentive compatible (i.e., not a good business decision) or not feasible. In a few cases the degrading was not deemed to be significant enough to greatly impact the market. In those cases we kept, there was some commitment regarding interoperation that was required for the merger to be allowed.
2.3 Identifying Relevant Technologies
With these fourteen cases in hand, we next identified the technology or technologies that were the subject of (potential) interoperability mandates. The cases often discussed many products and not all were necessarily related to interoperability mandates. Some cases discussed multiple distinct technologies subject to interoperability mandates. Some distinct cases discussed the same type of technology. We define the relevant technologies, within a case context, to be the technologies subject to interoperability mandates. Table IV (Appendix) lists relevant technologies for each case.
In some cases, we looked up mentions within the selected cases, of apparently similar cases about similar technologies. This led to detailed review of one EU case beyond the 14 selected (on WhatsApp and Messenger [^8] DMA.100024 Meta – number-independent interpersonal communications services).
2.4 Identifying SvI Considerations
Next, we collected discussion about the security implications of interoperability mandates—that is, SvI considerations—in each of our selected cases, across three sources: (1) the relevant company’s public-facing statements about the case, (2) mentions of security in relevant legal and regulatory proceedings, and (3) developer-facing information regarding the technical details of the relevant interoperation. By SvI considerations, we mean any actual, potential, perceived, or claimed tensions between security goals and mandated interoperation.
We collected such discussion by searching the press releases and developer documentation of the companies involved during the relevant time period, as well as Google searches for the company name and relevant technology alongside various combinations of the search terms ‘merger,’ ‘antitrust,’ ‘DMA,’ ‘EU,’ ‘Commission,’ ‘security,’ and ‘privacy.’ We note that this process might not have captured all company communications, especially for the older cases, as search interfaces and archives for press releases and developer documentation were variable. [^9] Our oldest case was decided in 2004, and the oldest relevant press release we were able to find is now only available on the Internet Archive. We were able to find company publications or news articles with official statements from companies about every case, although many did not mention security or mandated interoperability.
For mergers in particular, there was generally no discussion of the details of the commitments in public-facing venues. The only exception to this was the Google/Fitbit merger. As mergers are voluntary, they tend to involve situations in which companies are incentivized to make the merger appear favorable to shareholders and customers. In contrast to other antitrust proceedings, therefore, companies may be disincentivized from publicly suggesting that the terms of a merger will be disadvantageous to their security. [^10] Companies may make security arguments to the Commission that are not public [50], to which we did not have access.
Not every selected case involved an SvI concern. For some cases, we did not find any mentions of security at all: for example, the Cisco/Tandberg merger did not discuss the security of the underlying technology at all. [^11] Some of the mergers had surrounding discussion that said the merger and resulting interoperability (independent of any mandates) would improve consumer security as in the Intel/McAfee case.
Note
For those cases where we found no relevant security issues, we found a natural sub-categorization emerged, according to the type of interoperability mandate. All such cases we examined either required (1) that interoperability not be degraded in future, (2) that information related to interoperability be disclosed (but no further facilitation of interoperability was required), or (3) both. Table IV (Appendix) marks such cases as having ‘No’ SvI concern, followed by a sub-category (‘Degrade,’ ‘Disclose,’ or ‘Both’).
2.5 Systematization and Analysis
Our detailed review of these collected cases and materials, found three clear patterns—or categories—in the nature of the security arguments and the nature of changes needed to comply with the relevant interoperability mandate that form the basis for our taxonomy of SvI concerns (§3). The interplay of security and economic considerations that we observed repeatedly in the cases and materials formed the basis for our analysis questions (§4). We chose one to three case studies to feature for each taxonomy category, aiming to capture diverse and well-documented issues (§4).
3 Our Taxonomy